# Model Human Learners: Computational Models to Guide Instructional Design

### v1

@@ Introduction
In their seminal paper, ? (?) propose the concept of a Model Human Processor, a unified information-processing model that codifies psychological theories to predict human performance in user interfaces. They argue that interface designers could use this kind of model to make more informed design decisions that improve usability. In this paper, I propose an analogous concept: the Model Human Learner, a unified computational model that codifies cognitive and learning theories to predict human learning in instructional systems. I argue that instructional designers could use this kind of model to make more informed instructional design decisions that improve pedagogical effectiveness.
An analysis by ? (?) suggests that instructional designers face a crisis of choice—design spaces are combinatorial, contain trillions of possible configurations. How can instructional designer determine the best choices? Currently, they rely on A/B experiments to navigate this vast space, but human studies are costly and time consuming. Moreover, each A/B experiment provides only a single bit of information, meaning instructional designers are essentially playing a losing game of 20 questions with nature (?, ?).
@@
Prior research has explored several applications of computational models of learning. For example, ? (?) investigated their use for discovering cognitive models, while ? (?) explored how they can promote “learning-by-teaching,” where students learning by instructing simulated students. ? (?) and ? (?) examined their applications for tutor authoring, letting end users build tutors via tutoring rather than programming. Lastly, ? (?) and ? (?) studied their utility for theory testing.
@@ Discussion
I used the same simulation procedure as study 1, creating an agent for each student and connecting it to a machine-readable version of the box and arrows tutor, shown in Figure 4 to simulate learning. An analysis of the human data showed that the majority of human students solved the easy problems correctly on the first attempt. In contrast, only a few students got the hard problems correct on the first try because they are specifically designed to prevent prior knowledge use. Thus, for my analysis, I trained the agents on all the problems, but only analyzed their performance on the hard problems, because my model does not take into account prior knowledge, which seems to plays a major role in solving easy problems. [^4] Half of the simulated agents received prior training on 16 easy problems; I pre-trained the agents in case there was positive or negative transfer from the easy problems to the hard ones. The focus on hard problems makes it possible to compare agents and humans in a situation where they have matched prior knowledge. This subset of the human data contained 202 students solving 16 hard problems, for a total of 3,232 observations.

### v2

@@ Introduction
In their seminal paper, ? (?) propose the concept of a Model Human Processor, a unified information-processing model that codifies psychological theories to predict human performance in user interfaces. They argue that interface designers could use this kind of model to make more informed design decisions that improve usability. In this paper, I propose an analogous concept: a Model Human Learner, a unified computational model that codifies cognitive and learning theories to predict human learning in instructional systems. I argue that instructional designers could use this kind of model to make more informed instructional design decisions that improve pedagogical effectiveness.
An analysis by ? (?) suggests that instructional designers face a crisis of choice—design spaces are combinatorial, containing trillions of possible configurations. How can instructional designer determine the best choices? Currently, they rely on A/B experiments to navigate this vast space, but human studies are costly and time consuming. Moreover, each A/B experiment provides only a single bit of information, meaning instructional designers are essentially playing a losing game of 20 questions with nature (?, ?).
@@
Prior research has explored several applications of computational models of learning. For example, ? (?) investigated their use for discovering cognitive models, while ? (?) explored how they can promote “learning-by-teaching,” where students learn by instructing simulated students. ? (?) and ? (?) examined their applications for tutor authoring, letting end users build tutors via tutoring rather than programming. Lastly, ? (?) and ? (?) studied their utility for theory testing.
@@ Discussion
I used the same simulation procedure as study 1, creating an agent for each student and connecting it to a machine-readable version of the box and arrows tutor, shown in Figure 4 to simulate learning. An analysis of the human data showed that the majority of human students solved the easy problems correctly on the first attempt. In contrast, only a few students got the hard problems correct on the first try because they are specifically designed to prevent prior knowledge use. Thus, for my analysis, I trained the agents on all the problems, but only analyzed their performance on the hard problems, because my model does not take into account prior knowledge, which seems to play a major role in solving easy problems. [^4] Half of the simulated agents received prior training on 16 easy problems; I pre-trained the agents in case there was positive or negative transfer from the easy problems to the hard ones. The focus on hard problems makes it possible to compare agents and humans in a situation where they have matched prior knowledge. This subset of the human data contained 202 students solving 16 hard problems, for a total of 3,232 observations.